OpenStudy (anonymous):

solve each equation 3x - 1/ 6 - x +3/2 = 3x + 4/ 3

OpenStudy (anonymous):

\[\frac{3x-1}{6} - \frac{x +3}{2} = \frac{3x+4}{3}\] \[3x-1-3x-9=6x+8\] \[6x=-18\] \[x=-3\]